1, if in an a.p ,mtm=ntn  then show that tm+n=0


2. if the pth term of an a.p is q and the qth term is p , then find its (p+q) th term.


3. prove that for all finite values of a and b


(a-b)², (a+b)²,a²+b² are in a.p.


4. in the arithmatic  sequence 3,7,11, 15............ and 2,5,8..........each continued to 100 terms ,find how many terms are identical.


 

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear sougata

Ist question is not clear

2nd question:


let first term is =a  and common difference =d

p =a+(q-1)d  ............1

q =a+(p-1)d  ..............2


1-2

p-q = (q-p)d

d =-1

put value of d in equation 1

a =p+q -1


now p+q th term  = a+ ( p+q -1)d

                         = p+q -1 + ( p+q -1)*(-1)

                        =0


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin